40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 25 16:56:44 2022, Paco, Configuration, BHD, IFO recovery - IMC alignment 
    Reply  Wed May 25 17:34:47 2022, yuta, Configuration, BHD, IFO recovery - IFO alignment Screenshot_2022-05-25_17-47-57.png
       Reply  Thu May 26 19:55:43 2022, yuta, Configuration, BHD, Oplevs centered, BHD DCPDs are now online elog_1Y2.JPGelog_BHD.JPGelog_box.JPGScreenshot_2022-05-26_17-37-27_IFOaligned_OplevCentered.pngScreenshot_2022-05-26_20-35-02.png
          Reply  Fri May 27 17:45:53 2022, yuta, Configuration, BHD, BHD camera installed, GRY aligned 
             Reply  Thu Jun 2 20:05:37 2022, yuta, Configuration, PSL, IMC input power recovered to 1W, some alignment works Before.JPGAfter.JPG
                Reply  Fri Jun 3 12:13:58 2022, Paco, Configuration, CDS, Fix RFM channels  SoGreen.png
                   Reply  Fri Jun 3 15:22:51 2022, yuta, Update, LSC, Both arms locked with POY/POX, IR beam centered on TMs with ASS IRBeamsOnTMs.JPGScreenshot_2022-06-03_15-03-51.png
                      Reply  Fri Jun 3 17:42:50 2022, yuta, Update, LSC, MICH locks with AS55_Q Screenshot_2022-06-03_17-41-55.png
                         Reply  Sun Jun 5 19:46:40 2022, Paco, Update, LSC, Fixed IMC Trans sum issue 
                         Reply  Mon Jun 6 13:35:11 2022, Paco, Update, LSC, First calibrated spectra of MICH at AS55 Q Screenshot_2022-06-06_17-30-16_MICHOLTF.pngCalibrated_MICH_ERR.pdf
                            Reply  Fri Jun 17 16:22:21 2022, yuta, Update, LSC, Actuator calibration of BS. ITMX, ITMY, updated MICH displacement spectra from c1cal 6x
                               Reply  Thu Jul 7 18:18:19 2022, yuta, Update, LSC, Actuator calibration of ETMX and ETMX Screenshot_2022-07-05_14-52-01_OLTF.pngScreenshot_2022-07-05_14-54-03_TF.pngScreenshot_2022-07-05_14-56-41_Ratio.png
                                  Reply  Thu Jul 7 18:22:12 2022, yuta, Update, LSC, Actuator calibration of MC2 using Yarm TF.pngMC2.png
                                     Reply  Fri Jul 8 16:18:35 2022, rana, Update, LSC, Actuator calibration of MC2 using Yarm 
                                     Reply  Mon Jul 18 17:07:12 2022, yuta, Update, LSC, x4.12 added to ETMX coil outputs to balance with ETMY 
          Reply  Thu Jun 16 18:23:15 2022, Paco, Configuration, BHD, Recovering LO beam in BHD DCPDs Screenshot_2022-06-16_18-29-14_BHDLObeamISBACK.png
          Reply  Tue Jun 21 14:17:50 2022, yuta, Configuration, BHD, BHD DCPDs re-routed to c1sus2 C1X07ADC1.JPGBHDDCPDs.JPG
             Reply  Tue Jul 19 16:00:34 2022, yuta, Configuration, BHD, Fast channels for BHD DCPDs now available in c1lsc but not in c1hpc Screenshot_2022-07-19_14-26-39_c1hpc.pngScreenshot_2022-07-19_14-24-49_c1lsc.pngScreenshot_2022-07-19_15-51-25_GreenGreen.png
                Reply  Thu Jul 21 21:50:47 2022, Tega, Configuration, BHD, c1sus2 IPC update 
                   Reply  Fri Jul 22 15:05:26 2022, Tega, Configuration, BHD, c1sus2 shared memory and ADC fix Screenshot_2022-07-22_15-04-33_DCPD.pngScreenshot_2022-07-22_15-12-19_models.pngScreenshot_2022-07-22_15-15-11_ERR.pngScreenshot_2022-07-22_15-32-19_GDS.png
                      Reply  Fri Jul 22 17:46:10 2022, yuta, Configuration, BHD, c1sus2 watchdog update and DCPD ERR channels Screenshot_2022-07-22_17-48-25.png
                         Reply  Mon Jul 25 17:58:10 2022, Tega, Configuration, BHD, c1sus2 IPC dolphin issue update Screen_Shot_2022-07-25_at_5.43.28_PM.pngScreen_Shot_2022-07-25_at_5.21.10_PM.pngScreen_Shot_2022-07-25_at_5.30.58_PM.pngScreen_Shot_2022-07-25_at_5.35.19_PM.png
                            Reply  Mon Aug 1 18:42:39 2022, Tega, Configuration, BHD, c1sus2 IPC dolphin issue update c1sus2_dolphin.pngfb1_dxamin_status.pngdolphin_num_mem_init2.png
                               Reply  Tue Aug 2 17:25:18 2022, Tega, Configuration, BHD, c1sus2 dolphin IPC issue solved c1lsc_IPC_status.pngFE_lsmod_dependencies_c1sus2_b4_after_iop_unpdate.png
                      Reply  Mon Jul 25 18:09:41 2022, Tega, Configuration, BHD, BHD Homodyne Phase control MEDM screen Screen_Shot_2022-07-25_at_6.12.08_PM.pngScreen_Shot_2022-07-25_at_6.18.09_PM.png
Message ID: 16978     Entry time: Thu Jul 7 18:22:12 2022     In reply to: 16977     Reply to this: 16981   17014
Author: yuta 
Type: Update 
Category: LSC 
Subject: Actuator calibration of MC2 using Yarm 

(This is also a restore of elog 40m/16971 from Jul 5, 2022 at 17:36)

MC2 actuator calibration was also done using Yarm in the same way as we did in 40m/16970 (now 40m/16977).
The result is the following;
MC2 : -14.17e-9 /f^2 m/counts in arm length (-2.9905 times ITMY)
MC2 :   5.06e-9 /f^2 m/counts in IMC length
MC2 :  1.06e+05 /f^2 Hz/counts in IR laser frequency

What we did:
- Measured TF from C1:LSC-MC2_EXC to C1:LSC-YARM_IN1 during YARM lock using ETMY (see Attachment #1). Note that the sign of MC2 actuation and ITMY actuation is flipped.
- Took the ratio between ITM actuation and MC2 actuation to calculate MC2 actuation. For ITM actuation, we used the value measured using MICH (see 40m/16929). The average of the ratio in the frequency range 70-150 Hz was used (see Attachment #2).
- The actuation efficiency in meters in arm length was converted into meters in IMC length by multiplying it by IMCLength/ArmLength, where IMCLength=13.5 m is half of IMC round-trip length, ArmLength=37.79 m is the arm length.
- The actuation efficiency in meters in arm length was converted into Hz in IR laser frequency by multiplying it by LaserFreq/ArmLength, where LaserFreq=1064 nm / c is the laser frequency.

Files:
- Measurement files live in https://git.ligo.org/40m/measurements/-/tree/main/LSC/YARM
- Script for calculation lives at https://git.ligo.org/40m/scripts/-/blob/main/CAL/ARM/ETMActuatorCalibration.ipynb

Summary of actuation calibration so far:
BS   : 26.08e-9 /f^2 m/counts (see 40m/16929)
ITMX :  5.29e-9 /f^2 m/counts (see
40m/16929)
ITMY :  4.74e-9 /f^2 m/counts (see
40m/16929)
ETMX :  2.65e-9 /f^2 m/counts (0.5007 times ITMX)
ETMY : 10.91e-9 /f^2 m/counts (2.3017 times ITMY)

MC2 : -14.17e-9 /f^2 m/counts in arm length (-2.9905 times ITMY)
MC2 :   5.06e-9 /f^2 m/counts in IMC length

 

NOTE ADDED by YM on July 7, 2022

To account for the gain imbalance in ETMX, ETMY, MC2, LSC violin filter gains were set to:
C1:LSC-ETMX_GAIN = 4.12
C1:LSC-MC2_GAIN = -0.77
This is a temporary solution to make ETMX and MC2 actuation efficiencies from LSC in terms of arm length to be the same as ETMY 10.91e-9 /f^2 m/counts.

I think it is better to make C1:LSC-ETMX_GAIN = 1, and put 4.12 in C1:SUS-ETMX_TO_COIL gains. We need to adjust local damping gains and XARM ASS afterwards.
As for MC2, it is better to put -0.77 in LSC output matrix, since this balancing depends on LSC topology.

Attachment 1: TF.png  26 kB  Uploaded Thu Jul 7 19:37:19 2022  | Hide | Hide all
TF.png
Attachment 2: MC2.png  17 kB  Uploaded Thu Jul 7 19:37:31 2022  | Hide | Hide all
MC2.png
ELOG V3.1.3-